# Efficiently Ranking Software Variants with Minimal Benchmarks

### v1

@@ 1. Introduction
Benchmarking is a fundamental practice in software engineering that focuses on evaluating the qualities and performance of software systems. It is also essential in computer science for tasks such as optimization of algorithms or evaluation of system performance (Bartz-Beielstein et al., 2020; Olson et al., 2017; Kerschke et al., 2019; Malan and Engelbrecht, 2013; Muñoz Acosta et al., 2015; Moreau et al., 2022). In computational science, scientists use benchmarking to explore, compare and validate the performance of complex variants of simulations and models, facilitating advances in research and development (Donoho, 2024; Guyon and Viegas, 2020). Benchmarks can be seen as a special kind of test that is used to measure software performance such as execution time, energy consumption, or accuracy. Unlike unit or integration tests that return binary or ternary test results (i.e., Pass/Fail/inconclusive), benchmarking returns numerical evaluations, which makes them more challenging to work with. This is why software system benchmarking usually involves comparisons with other software variants. Hence, benchmarks are executed on different variants, whether they originate from different competing, alternative systems, or various configurations of the same system.
@@
Benchmarks and performance comparisons are crucial for understanding the performance characteristics of software variants, configuring and fine-tuning software, detecting regressions, and guiding the design of new systems. However, executing comprehensive benchmarks to achieve a complete understanding of software variants can be extremely costly in terms of computational resources and time. For instance, LLM leaderboards (Liu et al., 2023a) or SAT competitions (Froleyks et al., 2021; Heule et al., 2019) are two well-known examples of benchmarking competitions that require extensive computational resources to evaluate the performance of variants. The need for more efficient methods to maintain accurate rankings of software variants while minimizing the computational load is evident in several concrete scenarios:
@@
  • Performance modeling, extrapolation, and configuration tuning: Engineers often use mathematical models and simulations to predict software performance under various conditions or to find optimal configurations (Balsamo et al., 2004). By selecting a subset of representative benchmarks, they can efficiently train models that capture essential performance distributions. When benchmarking is training (Bartz-Beielstein et al., 2020), this process reduces the need for exhaustive execution, allowing engineers to efficiently identify performance bottlenecks and optimize system configurations based on workload sensitivity (e.g., as recently shown in (Mühlbauer et al., 2023; Lesoil et al., 2023)).
@@
However, all of these scenarios also highlight the potential dangers of reduced benchmarks that do not fully represent the entire set of benchmarks. To our knowledge, no previous work has addressed the problem of minimizing benchmarks while preserving the ranking of multiple software variants (see also more details in section 2). Previous research has extensively explored benchmarking and testing (Yoo and Harman, 2012), focusing on optimization, prioritization, and reduction of the test suite, but primarily for single systems, functional tests, and not for competing software variants. The works on software product lines (SPL) focus on the analysis of multiple products (Lima et al., 2020; Al-Hajjaji et al., 2017; Elbaum et al., 2014; Ensan et al., 2011; Kumar and Rajkumar, 2016; Lachmann et al., 2017, 2015; Lity et al., 2019; Lochau et al., 2012; Marijan et al., 2019, 2013; Marijan and Liaaen, 2017; Marijan et al., 2017; Prado Lima and Vergilio, 2020; Runeson and Engström, 2012; Wang et al., 2015, 2016; Xu et al., 2013), but assume a single code base and seek to minimize the number of test cases while usually maximizing the functional coverage of features or code (not performance).
Our optimization goal thus differs and consists in preserving the performance ranking of potentially independent software variants. In this paper, we propose a novel method to reduce the computational cost of benchmarks while still effectively ranking software variants using test suite optimization techniques. Our method, called BISS for BISection Sampling, is based on the fact that some tests cannot be removed and form the essential core of the benchmark. Furthermore, once some tests are removed, other tests become crucial as well. This enables us to use a bisection sampling routine to sample among relevant tests cleverly. We couple it with a divide-and-conquer algorithm to reduce the variance and speed up the search, and with an automatic restart (iterative solving) as long as we find improvements. Our contributions include:
@@
The remainder of this paper is organized as follows: section 2 discusses related work further. In section 3 we introduce the problem. In section 4 we first describe our method and then in section 5 propose baselines to compare to, including implementation details. We provide experimental results in section 6. Then we focus on selected case studies in  section 7. We discuss threats to validity in section 8. We discuss results and provide recommendations for developers and benchmark designers and users in section 9. Finally, we draw general conclusions and discuss future work in section 10.
@@ 2. Related work
Numerous work on regression testing and test suite minimization, selection, and prioritization has been published (Yoo and Harman, 2012). The principle is to improve the test suite of a single software system, typically to reduce computational cost as test suites tend to grow in size. Test suite minimization seeks to prune test suites from redundant test cases to reduce the number of tests to run. We pursue a related, yet different, goal with two key differences: we are interested in performance of software and we consider multiple software variants.
@@
Research in software performance engineering explores benchmarking at different levels (e.g., system-level tests and method/statement-level tests) and with different approaches (e.g., blackbox vs whitebox) (Laaber et al., 2020; Weyuker and Vokolos, 2000; Jiang and Hassan, 2015; Menascé, 2002; Blot and Petke, 2025; Nguyen et al., 2014; AlGhamdi et al., 2020). For example, there are techniques for microbenchmarking that measure the performance of e.g., single functions (Laaber et al., 2020; Maricq et al., 2018; Laaber et al., 2020; Costa et al., 2019). Ding et al. (Ding et al., 2020) studied how unit tests can be used to assess performance properties. All of these works do not aim to preserve the performance ranking of multiple variants.
@@
In close proximity to our work are contributions related to software product lines (SPLs) and configurable systems (Thüm et al., 2014; Rhein et al., 2018; Lima et al., 2020; Al-Hajjaji et al., 2017; Elbaum et al., 2014; Ensan et al., 2011; Kumar and Rajkumar, 2016; Lachmann et al., 2017, 2015; Lity et al., 2019; Lochau et al., 2012; Marijan et al., 2019, 2013; Marijan and Liaaen, 2017; Marijan et al., 2017; Prado Lima and Vergilio, 2020; Runeson and Engström, 2012; Wang et al., 2015, 2016; Xu et al., 2013; Chen et al., 2019). In this line of research, the goal is to consider a set of related products or configurations coming from a common code base (an SPL or a configurable system) and to optimize testing activities for these products. Existing SPL works explore the three traditional regression testing techniques for single software systems (Yoo and Harman, 2012), namely minimization (Wang et al., 2015), selection (Lity et al., 2019; Lochau et al., 2012; Wang et al., 2016; Xu et al., 2013), and prioritization (Al-Hajjaji et al., 2017; Lachmann et al., 2017, 2015) of test cases. Wang et al. propose to optimize test suites considering effectiveness measures such as test minimization percentage, pairwise coverage, fault detection capability (Wang et al., 2015) but these works ignore software performance properties and ranking information. Mendonça et al. introduce FeaTestSelPrio a feature-oriented test case selection and prioritization approach that improves failure detection and reduces test execution time by linking test cases to feature implementations and prioritizing based on feature coverage (Mendonça et al., 2024). Our work differs from several perspectives: (1) we target variants from multiple code bases; (2) we leave apart variant features: such features can be unavailable in our context, where we only have access to the software variants themselves; (3) performance calls to consider quantitative properties of software, whereas test suites are typically qualitative (e.g., Fail/Pass or Inconclusive); (4) our optimization goal is to preserve the (performance) ranking of software variants, while prior works consider failures or feature/code coverage.
@@
Instead of reducing benchmarks, a more common problem is the selection of relevant software configurations for a given problem. Several configuration sampling strategies have been proposed in the literature on SPLs and configurable systems, either to find and cover as many faults as possible (Han and Yu, 2016; Thüm et al., 2014), or to measure and learn the performance of the configuration space (Alves Pereira et al., 2021, 2020; Kaltenecker et al., 2020). Their focus is on the selection of configurations (variants) to test, rather than on the selection of test cases to run. Recent empirical findings from Mühlbauer et al. (Mühlbauer et al., 2023) and Lesoil et al. (Lesoil et al., 2023) show that program inputs (or workloads), typically part of a benchmark, and configuration options can result in substantial performance variations. The sheer volume of configurations and benchmark elements creates high computational costs, making it impractical to learn a performance model for each test case. These performance variations can alter the ranking of software variants, highlighting the need for careful benchmark selection to ensure accurate assessments. Nair et al. (Nair et al., 2017) tackle configuration optimization through rank-based selection assuming a given workload and relying on potentially imprecise performance models. In contrast, our approach focuses on reducing the benchmark suite itself to preserve the ranking across software variants, rather than identifying optimal configurations. Existing work, such as Nair et al., is not applicable to benchmark reduction, as it focuses on optimizing configurations within a single workload and does not address the challenge of selecting benchmarks to maintain ranking consistency across multiple variants.
@@
Multimorphic testing evaluates the coverage of a test suite with respect to a quantitative property by generating multiple software variants to reveal performance differences (Temple et al., 2018). Assesses performance test suite effectiveness by identifying tests that can ”kill” underperforming configurations and calculating a dispersion score to measure coverage across variants. While we operate in a similar context, our focus differs fundamentally: instead of the dispersion score, we aim to ensure that the minimized benchmark set preserves the correct performance ranking among competing variants, which is essential for various scenarios described above.
@@
The automated algorithm selection problem is subject to intensive research (Kerschke et al., 2019; Hutter et al., 2011, 2014; Xu et al., 2008; Thornton et al., 2013) with many applications in artificial intelligence (such as SAT, CSP, QBF, ASP, or scheduling (Kerschke et al., 2019)). Given a computational problem, a set of algorithms, and a specific problem instance to be solved, the challenge is to determine which of the algorithms can be selected to perform best on that instance. For example, SATZilla uses machine learning to select the most effective algorithm from a portfolio of SAT solvers for a given SAT formula (Xu et al., 2008). In addition, determining which of two algorithms performs best on a benchmark, without running the whole benchmark, has been discussed in (Matricon et al., 2021). However, the focus is on two variants and the cost of running variants, not on benchmarks. A critical factor in automated algorithm selection is the set of problem instances (benchmarks) used for e.g., training models to facilitate optimal algorithm selection. Our work contributes to this area by providing a software engineering methodology that can be applied to various AI problems.
@@ 3. Problem Statement: Ranked Test Suite Minimization (RTSM)
Individual performance values are usually difficult to get and verify. For example, checking that a system runs exactly in 2.42 seconds, as it has been registered from previous executions or predicted by an ML model, is very unlikely. Comparison of this with other observations coming from other variants’ performance estimations is usually needed. Ranking performance estimations are a way to evaluate how a system operates and is widely used (Bartz-Beielstein et al., 2020; Lesoil et al., 2023).
@@
$r=Index(\operatorname*{\arg\!\max}_{v\in\mathcal{A}}(perf(v,\mathcal{T}))$
The variant with the maximal performance estimation is ranked $1$ and so on. Note that this ranking computes a total ordering over the variants in $\mathcal{A}$.
@@
Note that test suite minimization is an instance of RTSM when one just ignores the ranking. As test suite minimization is NP-hard (Gotlieb and Marijan, 2014), we can deduce from the previous remark that RTSM is also NP-hard. In practice, for large instances, we will accept near-optimal solutions obtained by using heuristics, anytime algorithms, or probabilistic approaches.
@@
Unfortunately, there is no ground truth; therefore, we will compare the rankings obtained using the solution to the original solution. There are multiple metrics to compare the rankings, the two most commonly used ones are Kendall’s (Kendall, 1938) and Spearman’s (Spearman, 1904) rank correlation coefficient. Both give a score between -1, the two vectors offer opposite rankings, and 1, the two vectors offer exactly the same rankings. We chose to use Kendall’s, which is directly proportional to the number of variants correctly ranked. Choosing Spearman or Kendall does not matter when the aim is a score of 1, that is, the same ranking, because they both coincide for this data point. Both tests are non-parametric hypothesis tests on the relations between the two rankings. Kendall can be easily computed, assume that a pair of variants are concordant if they are ranked in the same order in both rankings and discordant otherwise, then we have:
@@ 4. Benchmark minimization with BISS
As explained in section 2, most existing methods for minimizing the test suite require prior information such as test results, historical data, etc. However, the scenarios for RTSM are very diverse, requiring handcrafted features by experts for their domain if we want to use a white-box approach, which are often not available or are too costly to acquire. Instead, we choose to look at black-box approaches. We have no features about the variants; we have just an identifier to be able to identify them, and this is the same for the tests. So, as input, we only have the performance matrix, that is, a matrix that contains the performance of each variant on each test. Then techniques such as principal component analysis (PCA) or racing techniques (López-Ibáñez et al., 2016) should solve this problem. We will see that we tested these techniques and they do not solve our problem. Indeed, these classic techniques can work to some extent, but we are interested in objectives that are a loose proxy of our interest: the ranking of variants. Whereas most of these techniques are interested in more fine-grained information: the magnitude of the differences, they care if you are twice as fast or ten times as fast, whereas in RTSM it does not matter.
@@
The core idea is that we do not care about the individual performances of the tests; we only care about the relative performance of a variant to another across tests.
Definition 3.2 gives us the possibility of assigning weights to $W$: Given a subset of the test set, we can use a linear regression to optimize $W$ to obtain the best coefficients. Let $P\in\mathbb{R}^{|Variants\times|\mathcal{T}|}$ be the performance matrix of the variants in the test, then, given $\mathcal{T}$’ of $\mathcal{T}$, we can optimize:
$\forall i\in\text{Variants, }P_{i\mathcal{T}^{\prime}}W^{T}=\sum_{j=0}^{|\mathcal{T}|}P_{ij}$
This is a linear regression problem where $W$ is a vector that assigns a weight to each test in $\mathcal{T}^{\prime}$, which is the set of parameters that we try to find for each variant, given their performance in the variants, to predict their performance on all tests. Our goal is therefore to sample relevant subsets $\mathcal{T}^{\prime}$. Formally, we assume that we have the procedure $solves_{T}(T^{\prime})$ which solves the linear regression and checks if the Kendall target is reached with $T^{\prime}$ as a solution to the problem with the complete test set being $T$, if it is, then the procedure returns truetrue otherwise, it returns falsefalse.
As we have explained earlier, the problem is NP-hard, and covering all possible solutions is intractable for small test suites; for a test suite of size 30, there are more than a billion possible solutions. The process is a bit akin to dimensionality reduction; each test acts as a dimension, and we want to reduce the dimension while keeping the ranks intact. Taking inspiration from the Johnson–Lindenstrauss lemma, which states that a random projection in the ”low” dimension has a high probability of keeping the original distances between the data points in the original high dimension. The dimension that the lemma gives us is dependent logarithmically on the original dimension and grows inversely proportional to the square of the minimum distance between two data points. In other words, the closer two variants have similar performance, the more likely it is that we are going to need more tests to rank them correctly. Therefore, our objective is to sample subsets of tests intelligently that are likely to result in a solution.

### v2

@@ 1. Introduction
Benchmarking is a fundamental practice in software engineering that focuses on evaluating the qualities and performance of software systems. It is also essential in computer science for tasks such as optimization of algorithms or evaluation of system performance (bartzbeielstein2020benchmarkingoptimizationbestpractice; Olson2017; Kerschke2019; Malan2013; Munoz2015; benchopt). In computational science, scientists use benchmarking to explore, compare and validate the performance of complex variants of simulations and models, facilitating advances in research and development (Donoho2024Data; Guyon2020). Benchmarks can be seen as a special kind of test that is used to measure software performance such as execution time, energy consumption, or accuracy. Unlike unit or integration tests that return binary or ternary test results (i.e., Pass/Fail/inconclusive), benchmarking returns numerical evaluations, which makes them more challenging to work with. This is why software system benchmarking usually involves comparisons with other software variants. Hence, benchmarks are executed on different variants, whether they originate from different competing, alternative systems, or various configurations of the same system.
@@
Benchmarks and performance comparisons are crucial for understanding the performance characteristics of software variants, configuring and fine-tuning software, detecting regressions, and guiding the design of new systems. However, executing comprehensive benchmarks to achieve a complete understanding of software variants can be extremely costly in terms of computational resources and time. For instance, LLM leaderboards (liu2023codegeneratedchatgptreally) or SAT competitions (froleyks2021sat; heule2019sat) are two well-known examples of benchmarking competitions that require extensive computational resources to evaluate the performance of variants. The need for more efficient methods to maintain accurate rankings of software variants while minimizing the computational load is evident in several concrete scenarios:
@@
  • Performance modeling, extrapolation, and configuration tuning: Engineers often use mathematical models and simulations to predict software performance under various conditions or to find optimal configurations (balsamo04modelbased). By selecting a subset of representative benchmarks, they can efficiently train models that capture essential performance distributions. When benchmarking is training (bartzbeielstein2020benchmarkingoptimizationbestpractice), this process reduces the need for exhaustive execution, allowing engineers to efficiently identify performance bottlenecks and optimize system configurations based on workload sensitivity (e.g., as recently shown in (muhlbauer2023; lesoil2023)).
@@
However, all of these scenarios also highlight the potential dangers of reduced benchmarks that do not fully represent the entire set of benchmarks. To our knowledge, no previous work has addressed the problem of minimizing benchmarks while preserving the ranking of multiple software variants (see also more details in section 2). Previous research has extensively explored benchmarking and testing (harman2012RT), focusing on optimization, prioritization, and reduction of the test suite, but primarily for single systems, functional tests, and not for competing software variants. The works on software product lines (SPL) focus on the analysis of multiple products (10.1145/3382025.3414967; Al-Hajjaji2017; Elbaum2014; Ensan2011; Kumar2016; Lachmann2017; Lachmann2015; Lity2019; Lochau2012; Marijan2019; Marijan2013; Marijan2017; Marijan2017titan; PradoLima2020; Runeson2012; wang_cost-effective_2015; Wang2016; Xu2013), but assume a single code base and seek to minimize the number of test cases while usually maximizing the functional coverage of features or code (not performance).
Our optimization goal thus differs and consists in preserving the performance ranking of potentially independent software variants. In this paper, we propose a novel method to reduce the computational cost of benchmarks while still effectively ranking software variants using test suite optimization techniques. Drawing from the fact that the problem is NP-hard, we turn to heuristics with anytime algorithms because even small reductions can provide benefits; there is no need for exact optimality most of the time. Our method, called BISS for BISection Sampling, is based on the fact that some tests cannot be removed and form the essential core of the benchmark. Furthermore, once some tests are removed, other tests become crucial as well. This enables us to use a bisection sampling routine to sample among relevant tests cleverly. We couple it with a divide-and-conquer algorithm to reduce the variance and speed up the search, and with an automatic restart (iterative solving) as long as we find improvements. Our contributions include:
@@
The remainder of this paper is organized as follows: section 2 discusses related work further. In section 3 we introduce the problem. In section 4 we first describe our method and then in section 5 propose baselines to compare to, including implementation details. We provide experimental results in section 6. Then we focus on selected case studies in  section 7. We discuss threats to validity in section 8. We draw conclusions, provide recommendations for developers and benchmark designers and users, and discuss future work in section 9.
@@ 2. Related work
Numerous work on regression testing and test suite minimization, selection, and prioritization has been published (harman2012RT). The principle is to improve the test suite of a single software system, typically to reduce computational cost as test suites tend to grow in size. Test suite minimization seeks to prune test suites from redundant test cases to reduce the number of tests to run. We pursue a related, yet different, goal with two key differences: we are interested in performance of software and we consider multiple software variants.
@@
Research in software performance engineering explores benchmarking at different levels (e.g., system-level tests and method/statement-level tests) and with different approaches (e.g., blackbox vs whitebox) (microbenchmarks2020; weyuker2000experience; jian2015survey; menasce2002load; blot2022comprehensivesurveybenchmarksautomated; nguyen2014industrial; alghamdi2020towards). For example, there are techniques for microbenchmarking that measure the performance of e.g., single functions (microbenchmarks2020; Maricq2018; microbenchmarks2020; costa2019whats). Ding et al. (ding2020towards) studied how unit tests can be used to assess performance properties. All of these works do not aim to preserve the performance ranking of multiple variants.
@@
In close proximity to our work are contributions related to software product lines (SPLs) and configurable systems (thum2014classification; rhein2018variability; 10.1145/3382025.3414967; Al-Hajjaji2017; Elbaum2014; Ensan2011; Kumar2016; Lachmann2017; Lachmann2015; Lity2019; Lochau2012; Marijan2019; Marijan2013; Marijan2017; Marijan2017titan; PradoLima2020; Runeson2012; wang_cost-effective_2015; Wang2016; Xu2013; sway). In this line of research, the goal is to consider a set of related products or configurations coming from a common code base (an SPL or a configurable system) and to optimize testing activities for these products. Existing SPL works explore the three traditional regression testing techniques for single software systems (harman2012RT), namely minimization (wang_cost-effective_2015), selection (Lity2019; Lochau2012; Wang2016; Xu2013), and prioritization (Al-Hajjaji2017; Lachmann2017; Lachmann2015) of test cases. Wang et al. propose to optimize test suites considering effectiveness measures such as test minimization percentage, pairwise coverage, fault detection capability (wang_cost-effective_2015) but these works ignore software performance properties and ranking information. Mendonça et al. introduce FeaTestSelPrio a feature-oriented test case selection and prioritization approach that improves failure detection and reduces test execution time by linking test cases to feature implementations and prioritizing based on feature coverage (mendonca_feature-oriented_2024). Our work differs from several perspectives: (1) we target variants from multiple code bases; (2) we leave apart variant features: such features can be unavailable in our context, where we only have access to the software variants themselves; (3) performance calls to consider quantitative properties of software, whereas test suites are typically qualitative (e.g., Fail/Pass or Inconclusive); (4) our optimization goal is to preserve the (performance) ranking of software variants, while prior works consider failures or feature/code coverage.
@@
Instead of reducing benchmarks, a more common problem is the selection of relevant software configurations for a given problem. Several configuration sampling strategies have been proposed in the literature on SPLs and configurable systems, either to find and cover as many faults as possible (han2016empirical; thum2014classification), or to measure and learn the performance of the configuration space (alvespereira:hal-02148791; alvespereira:hal-02356290; kaltenecker2020interplay). Their focus is on the selection of configurations (variants) to test, rather than on the selection of test cases to run. Recent empirical findings from Mühlbauer et al. (muhlbauer2023) and Lesoil et al. (lesoil2023) show that program inputs (or workloads), typically part of a benchmark, and configuration options can result in substantial performance variations. The sheer volume of configurations and benchmark elements creates high computational costs, making it impractical to learn a performance model for each test case. These performance variations can alter the ranking of software variants, highlighting the need for careful benchmark selection to ensure accurate assessments. Nair et al. (nair2017) tackle configuration optimization through rank-based selection assuming a given workload and relying on potentially imprecise performance models. In contrast, our approach focuses on reducing the benchmark suite itself to preserve the ranking across software variants, rather than identifying optimal configurations. Existing work, such as Nair et al., is not applicable to benchmark reduction, as it focuses on optimizing configurations within a single workload and does not address the challenge of selecting benchmarks to maintain ranking consistency across multiple variants.
@@
Multimorphic testing evaluates the coverage of a test suite with respect to a quantitative property by generating multiple software variants to reveal performance differences (multimorphic). It assesses performance test suite effectiveness by identifying tests that can ”kill” under-performing configurations and calculating a dispersion score to measure coverage across variants. While we operate in a similar context, our focus differs fundamentally: instead of the dispersion score, we aim to ensure that the minimized benchmark set preserves the correct performance ranking among competing variants, which is essential for various scenarios described above.
@@
The automated algorithm selection problem is subject to intensive research (Kerschke2019; hutter2011; Hutter2014; xu2008; thornton2013) with many applications in artificial intelligence (such as SAT, CSP, QBF, ASP, or scheduling (Kerschke2019)). Given a computational problem, a set of algorithms, and a specific problem instance to be solved, the challenge is to determine which of the algorithms can be selected to perform best on that instance. For example, SATZilla uses machine learning to select the most effective algorithm from a portfolio of SAT solvers for a given SAT formula (xu2008). In addition, determining which of two algorithms performs best on a benchmark, without running the whole benchmark, has been discussed in (matricon21comparison). However, the focus is on two variants and the cost of running variants, not on benchmarks. A critical factor in automated algorithm selection is the set of problem instances (benchmarks) used for e.g., training models to facilitate optimal algorithm selection. Our work contributes to this area by providing a software engineering methodology that can be applied to various AI problems.
@@ 3. Problem Statement: Ranked Test Suite Minimization (RTSM)
Individual performance values are usually difficult to get and verify. For example, checking that a system runs exactly in 2.42 seconds, as it has been registered from previous executions or predicted by an ML model, is very unlikely. Comparison of this with other observations coming from other variants’ performance estimations is usually needed. Ranking performance estimations are a way to evaluate how a system operates and is widely used (bartzbeielstein2020benchmarkingoptimizationbestpractice; lesoil2023).
@@
$r(v)=|\{v^{\prime}\in\mathcal{A}\ |\ perf(v,\mathcal{T})<perf(v^{\prime},\mathcal{T})\}|$
The variant with the maximal performance estimation is ranked $1$ and so on. Note that this ranking computes a total ordering over the variants in $\mathcal{A}$. The above formula assumes that there is no tie. In practice, ties might occur and a tie-breaking strategy is required.
In the experiments of this paper, ties are broken by ranking tied solvers alphabetically but ties never occur.
@@
Note that test suite minimization is an instance of RTSM when one just ignores the ranking. As test suite minimization is NP-hard (goliebissta14), we can deduce from the previous remark that RTSM is also NP-hard. That means, while finding the minimal subset is NP-hard, finding a solution remains easy: $\mathcal{T}$ is a solution. In this work, we do not attempt to find an optimal solution to this NP-hard problem, but propose an empirically efficient heuristic solution, designed for practical applications.
@@
We will use the weighted approach whenever we want to reduce the cost of running a test suite and not just the number of tests. This is of course also NP-hard, that is finding the test suite with minimal cost is NP-hard.
@@
Unfortunately, there is no ground truth; therefore, we will compare the rankings obtained using the solution found and the original ranking prior to minimization. There are multiple metrics to compare the rankings, the two most commonly used ones are Kendall’s (kendall) and Spearman’s (spearman) rank correlation coefficient. Both give a score between -1, the two vectors offer opposite rankings, and 1, the two vectors offer exactly the same rankings. We chose to use Kendall’s, which is directly proportional to the number of variants correctly ranked. Choosing Spearman or Kendall does not matter when the aim is a score of 1, that is, the same ranking, because they both coincide for this data point. Both tests are non-parametric hypothesis tests on the relations between the two rankings. Kendall can be easily computed, assume that a pair of variants are concordant if they are ranked in the same order in both rankings and discordant otherwise, then we have:
@@ 4. Benchmark minimization with BISS
As explained in section 2, most existing methods for minimizing the test suite require prior information such as test results, historical data, etc. They are all mostly practical approaches offering no guarantees due to the scalability required. However, the scenarios for RTSM are very diverse, requiring handcrafted features by experts for their domain if we want to use a white-box approach, which are often not available or are too costly to acquire. Instead, we choose to look at black-box approaches. We have no features about the variants; we have just an identifier to be able to identify them, and this is the same for the tests. So, as input, we only have the performance matrix, that is, a matrix that contains the performance of each variant on each test. Then techniques such as principal component analysis (PCA) or racing techniques (LOPEZIBANEZ201643) should solve this problem. We will see that we tested these techniques and they do not solve our problem. Indeed, these classic techniques can work to some extent, but we are interested in objectives that are a loose proxy of our interest: the ranking of variants. Whereas most of these techniques are interested in more fine-grained information: the magnitude of the differences, they care if you are twice as fast or ten times as fast, whereas in RTSM it does not matter.
@@
We adopt an empirical approach in which the core idea is that we do not care about the individual performances of the tests; we only care about the relative performance of a variant to another across tests.
Definition 3.2 gives us the possibility of assigning weights to $W\in\mathbb{R}^{|\mathcal{T}^{\prime}|\times 1}$: Given a subset of the test set, we can use a linear regression to optimize $W$ to obtain the best coefficients. Let $P\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{T}|}$ be the performance matrix of the variants on the tests, then, given a subset $\mathcal{T}$’ of $\mathcal{T}$, we can optimize:
[…]
where $\mathbf{1}_{|\mathcal{T}|\times 1}\in\mathbb{R}^{|\mathcal{T}|\times 1}$ is a matrix of ones and $P_{*\mathcal{T}^{\prime}}$ is the matrix $P$ where only the columns concerning the tests in $\mathcal{T}^{\prime}$ were kept. This is an ordinary least squares where the goal is to predict the aggregated performance on the full set of tests $\mathcal{T}$ from the set of tests that were kept $\mathcal{T}^{\prime}$. Our goal is therefore to sample relevant subsets $\mathcal{T}^{\prime}$. Formally, we assume that we have the procedure $\text{findWeightsAndCheck}_{T}(T^{\prime})$, that solves this linear regression and checks if the Kendall target is reached with $T^{\prime}$ as a solution to the problem with the complete test set being $T$. If it is, then the procedure returns truetrue, otherwise it returns false. $\text{findWeightsAndCheck}_{K\cup P}(K)$ has complexity O($|\mathcal{A}||K|^{2}|P|)$, where $\mathcal{A}$ is the set of variants. It is polynomial, but this is expected, as it only solves a linear regression problem and then computes if the solution is valid within the target Kendall value. It does not find the candidate subset $K$ nor does it prove that the solution is optimal.
As we have explained earlier, the problem is NP-hard, and covering all possible solutions is intractable for small test suites; for a test suite of size 30, there are more than a billion possible solutions. This is why our approach does not focus on finding an optimal solution. Instead, the goal is to produce a solution within the constraints. Notice that providing a solution can be considered trivial since in all formulations the initial test suite $\mathcal{T}$ is a valid solution. Our approach focuses on sampling better candidates than $\mathcal{T}$. The process is a bit akin to dimensionality reduction; each test acts as a dimension, and we want to reduce the dimension while keeping the ranks intact. We will take inspiration from the Johnson–Lindenstrauss lemma (johnson1984extensions), which states that a random projection in the “low” dimension has a high probability of keeping the original distances between the data points in the original high dimension. The dimension that the lemma gives us is logarithmically dependent on the original dimension and grows inversely proportional to the square of the minimum distance between two data points. In other words, the more similar the performance of two variants is, the more likely it is that more tests are needed to rank them correctly. Therefore, our objective is to sample subsets of tests that are likely to result in a solution.